Sophia's Restaurant served 6,800 meals last quarter. Sophia recorded the following costs with those meals. Variable costs: Ingredients used $ 17,940 Direct labor 14,100 Indirect materials and supplies 8,900 Utilities 5,300 Fixed costs: Managers' salaries 31,000 Rent 27,000 Depreciation on equipment (straight-line, time basis) 11,000 Other fixed costs 12,090 Required: Unit variable costs and total fixed costs are expected to remain unchanged next quarter. Calculate the unit cost and the total cost if 6,120 meals are served next quarter. (Round "Unit costs" answer to 2 decimal places.)
Answer #1

Answer

--Total Cost of 6120 meals = $ 122,706

--Unit cost for 6120 meals = $ 20.05 per unit

--Workings

Total Cost Unit Cost
Variable cost:
Ingrediants $16,146 $2.64
Direct labor $12,690 $2.07
Indirect material & supplies $8,010 $1.31
Utilities $4,770 $0.78
Total Variable cost $41,616 $6.80
Fixed Cost:
Manager's salary $31,000 $5.07
Rent $27,000 $4.41
Depreciation on equipment $11,000 $1.80
Other fixed cost $12,090 $1.98
Total Fixed Cost $81,090 $13.25
Total Cost $122,706 $20.05

--Calculations

Total Cost

Unit Cost

Variable cost:

Ingrediants

=6120*17940/6800

=17940/6800

Direct labor

=6120*14100/6800

=14100/6800

Indirect material & supplies

=6120*8900/6800

=8900/6800

Utilities

=6120*5300/6800

=5300/6800

Fixed Cost:

Manager's salary

31000

=31000/6120

Rent

27000

=27000/6120

Depreciation on equipment

11000

=11000/6120

Other fixed cost

12090

=12090/6120